OpenStudy (anonymous):

The functions f and g are given by f(x) = x^2+2 and g(x)= x+1^ (1/2) determine f(f(x))

OpenStudy (anonymous):

i have no clue what they mean with this

Parth (parthkohli):

Are you sure it's f(f(x)) and not g(f(x)) or f(g(x))?

OpenStudy (anonymous):

f(g(x)) aswell , g(f(x)) and f(f(x-1))

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(f(x)) = f(x^2 + 2) }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(x^2 + 2) = (x^2 + 2)^2 + 2 }\)

OpenStudy (anonymous):

a detailed answer to one of those would be nice i could use an example

OpenStudy (anonymous):

Willbert, it simply means you put the function inside itself.

OpenStudy (anonymous):

\(f(f(x))\) means "put" \(f(x)\) where \(x\) normally is.

Parth (parthkohli):

A function of a function :)

OpenStudy (anonymous):

ok ok I'll be backk later : ) I should go read the text book

OpenStudy (anonymous):

It always helps. I <3 my textbooks.

OpenStudy (anonymous):

What's the name of your math textbook @Limitless for algebra?

OpenStudy (anonymous):

Algebra (highschool level) is best covered via online texts. Purplemath.com is extremely thorough. Sometimes too thorough. Paul's online notes is good for more advanced things such as Precalc and Calculus. As for even higher subjects, SOSmath.com surprises me--they include things on differential equations and don't really stop there at all, lol. Khanacademy is also excellent. But, to be honest with you, I have probably done over 10,000 math problems in my entire life. That's what it comes down to at the end of the day: Dedication. Oh, and I highly recommend reading Wikipedia often. I can't tell you how many times I've read something I barely understood and then ended up using it later. :)

OpenStudy (anonymous):

Oh, cool thanks I was looking for a textbook to borrow from the library that has a lot of question that I could practice but I'll try those sites. Do you know where the practice questions are on Paul's online notes for algebra?

OpenStudy (anonymous):

Paul seems to include problems in the notes. Here is one example http://tutorial.math.lamar.edu/Classes/Alg/IntegerExponents.aspx.

OpenStudy (anonymous):

Yeah I noticed... but I saw practice questions somewhere and don't recall where...

OpenStudy (anonymous):

I'm not particularly familiar with where they are, either. Just google worksheets or practice problems for specific topics and you can find them. :D

OpenStudy (anonymous):

Okay :). Thanks. Do you know how to do the question I just posted?

OpenStudy (anonymous):

what question?